OpenStudy (anonymous):

Solve 5x + 2y = 4 for y.

OpenStudy (konradzuse):

Well, we want to isolate y, do you know how we should do that?

OpenStudy (hba):

1) Subtract both sides by 5x 2) divide both sides by 2

OpenStudy (anonymous):

@KonradZuse Well i know i have to substitute x with something else, i think..Right?, idk:(

OpenStudy (anonymous):

@hba, but its not 4x so i cant do that ..?

OpenStudy (hba):

You can do that but you cannot subtract them.

OpenStudy (hba):

@JessNicole Just follow me.

OpenStudy (hba):

Subtract both sides by 5x

OpenStudy (anonymous):

It'd be 2y = -1x ?

OpenStudy (hba):

No it would be 2y=4-5x You cannot subtract,You said Remember.

OpenStudy (hba):

Now divide both sides by 2

OpenStudy (anonymous):

y = -0.5x

OpenStudy (hba):

Whaaaaaat ?

OpenStudy (hba):

\[\frac{ 2y }{ 2}=\frac{ 4-5x }{ 2}\]

OpenStudy (anonymous):

4-5=-1...-1/2=-0.5x .. it doesnt work

OpenStudy (anonymous):

@hba

OpenStudy (hba):

Just keep it simple \[y=\frac{ 4-5x }{ 2 }\]

OpenStudy (hba):

@JessNicole Got it ?

OpenStudy (anonymous):

yea, some what

OpenStudy (hba):

Haha :P ok :)